OpenStudy (lifeisadangerousgame):

I have an exam tomorrow and I'm struggling with composite functions, particularly finding 2 to 3 simpler ones to describe one and ones with fractions.

OpenStudy (lifeisadangerousgame):

\[f(x) = \frac{ x + 2 }{ x - 3}; g(x) = \frac{ 1 }{ x + 1 }\] Find formulas for f(g(x)) and g(f(x))That's one, if you wouldn't mind showing me how to do it step by step I appreciate it.

OpenStudy (dumbcow):

its all about replacing the variable "x" with whatever the other function is ---> f(g) means that the variable is g(x) not "x" \[f(g(x)) = \frac{g(x) + 2}{g(x) - 3} = \frac{\frac{1}{x+1} +2}{\frac{1}{x+1} -3}\]

OpenStudy (dumbcow):

\[g(f(x)) = \frac{1}{f(x) +1} = \frac{1}{\frac{x+2}{x-3}+1}\]

OpenStudy (lifeisadangerousgame):

Okay, that makes sense, do you mind doing another one?

OpenStudy (dumbcow):

sure

OpenStudy (lifeisadangerousgame):

\[f(x) = \frac{ x + 2 }{ x - 3 }; g(x) = \frac{ 1 }{ x + 1 }\] I tried f(g(x)) and I got (2x + 2)/(-3x - 3) Is that right?

OpenStudy (lifeisadangerousgame):

Wait. I messed up, that's the same as the first problem

OpenStudy (dumbcow):

almost , should be (2x+3)/(-3x-2)

OpenStudy (lifeisadangerousgame):

Why's that?

OpenStudy (dumbcow):

\[\frac{1}{x+1}+\frac{2(x+1)}{x+1} = \frac{2x+3}{x+1}\]

OpenStudy (lifeisadangerousgame):

I'm still a bit lost, I'm sorry

OpenStudy (dumbcow):

hmm ok i was just combining terms into 1 fraction.... anyway you had a different question

OpenStudy (lifeisadangerousgame):

\[f(x) = \frac{ t-2 }{ t +3 }, g(x) = \frac{ 1 }{ (t+2)^2 }\]

OpenStudy (lifeisadangerousgame):

Actually, I'm still lost with how to do the first problem...I don't understand how it's )(2x + 3)/(-3x -2)

OpenStudy (mathmale):

Dumbcow did a good job of showing his work (above). It'd be easier for potential helpers to help you if you'd show (share) any work you've done yourself. By "first problem," do you mean your original post?